OpenStudy (anonymous):

describe how to make 100.0ml of a 0.10m solution of nah2po4 if a 0.50m stock solution is available.

OpenStudy (mos1635):

n2 moles in final solution C2=0.1 M V2=0.1 Lt n2=C2*V2=0.001 moles n2=n1 n2=C1*V1 V1 = ....